OpenStudy (anonymous):

Set up a proportion for the following situation.It takes Todd 8 minutes to eat 5 pies at a pie eating contest. How many pies can Todd eat in 15 minutes?

OpenStudy (nathan917):

I would say around 8 pieces

OpenStudy (anonymous):

I know it is 8, but I don't know how to set up a proportion

OpenStudy (anonymous):

How can you write that so that 5/8 = x/15?

OpenStudy (nathan917):

a proportion is set up to relate 2 relations. this is kinda confusing, but all i mean is create 2 fractions that will be equal to each other for example: 3 out of every 10 shirts that were sold in a day were blue. if 50 total shirts were sold, then how many were blue? you would set up proportions --> 2 fractions Fraction 1: 3/10 3 out of 10 were blue Fraction 2: X/50 How many out of 50 were blue set these equal to each other. 3/10=X/50 then you cross multiply 10X=3(50) Multiply 10X=150 Divide X=15 15 out of 50 shirts were blue

OpenStudy (nathan917):

That a good example to set up a proportion?

OpenStudy (anonymous):

yes it is